OpenStudy (anonymous):

suppose dice were rolled 248 times. Find the mean and standard deviation for a sum of seven?

OpenStudy (dumbcow):

\[mean = np = 248*\frac{1}{6}\] \[std = \sqrt{np(1-p)} = \sqrt{248*\frac{1}{6}*\frac{5}{6}}\]

OpenStudy (anonymous):

How many dice?